OpenStudy (anonymous):

5+3i/4-2i written as a complex number?

jhonyy9 (jhonyy9):

\[5+3i/4-2i = (5+3i)(4+2i)/(4-2i)(4+2i) =(5+3i)(4+2i)/(16-(-4))\]

jhonyy9 (jhonyy9):

(5+3i)(4+2i)/20 =(20+10i+12i-6)/20 =(14+22i)/20